OpenStudy (nani200):

Find a cubic function with the given zeros. 7, -3, 2

OpenStudy (mww):

There an an infinite number of cubic functions you can generate In general \[y = k(x - a)(x-b)(x-c)\] if a cubic with zeroes at x = a, b and c k is just a real, non zero constant (any number just not zero)

OpenStudy (nani200):

here are the answer choices A. x^3 - 6x^2 -13x - 42 B. x^3 - 6x^2 +13x +42 C. x^3 -6x^2 -13x + 42 D. x^3 + 6x^2 -13x +42

OpenStudy (mww):

well do as I said and expand the parentheses. use k = 1 to get a monic cubic.

OpenStudy (mww):

actually there is an easier way

OpenStudy (mww):

you can use sum, product of roots

OpenStudy (nani200):

i dont know how to expand them

OpenStudy (nani200):

how do you do that

OpenStudy (mww):

For monic cubic with roots 7,-3, and 2, you can write this as: y = (x - 7)(x+3)(x-2) Expand (x-7)(x+3)(x-2) = (x^2 +3x - 7x -21)(x-2) = (x^2 -4x -21)(x-2) = x^3 - 2x^2 - 4x^2 + 8x -21x +42 = X^3 - 6x^2 -13x + 42

OpenStudy (nani200):

thank you!

OpenStudy (nani200):

so its c